if A is 3 yrs older than B and B is twice of the age of C..the total sum of ages are 30 then what is the age of B?
Answer by Abbey(339)

A is 3 years older than B:
A=3+B
B is twice the age of C:
B=2C
so A compared to C is
A=3+2C
the sum of their ages is 30:
(3+2C)+(2C)+C=30
3+2C+2C+C=30
3+5C=30
subtract 3 from both sides:
5C=27
divide both sides by 5
C=27/5 or 5 and 2/5 years
so B is
2(27/5) =54/5 = 10 and 4/5 years
and A is 13 and 4/5 years
check:
10 and 4/5 + 13 and 4/5 + 5 and 2/5 = 30, so it is correct